OpenStudy (anonymous):

Use the following equations to solve the problem. final distance = initial distance + (speed × time) df = di + vt 1.00 mi = 1.61 km For an initial distance of 11.3 km, a speed of 35.0 miles per hour, and a time of 3.00 hours, what is the final distance in km?

OpenStudy (anonymous):

49.3 km 116 km 76.5 km 180 km

OpenStudy (shamim):

35 miles=? km

OpenStudy (anonymous):

you need to convert 35 miles per hour to km per hour... the conversion for miles to km is given already and @shamim is asking you to convert it...

OpenStudy (anonymous):

di=11.3km v=56.35km/hr t=3hr df=di+vt =11.3+(56.35*3) =180.35km